OpenStudy (anonymous):

Aldo works for a moving company after school for 4 hours each day on Monday through Thursday afternoons. On Saturday, he works for 6 hours. At $7.50 per hour, with no overtime, how much does he earn each week?

OpenStudy (anonymous):

4 hours x 4 days = 16 hours + 6 hours (Saturday) = 22 hours/week

OpenStudy (anonymous):

number of hours x pay = earnings so 22 x 7.50 = $165

OpenStudy (anonymous):

ooo Oakii THanks :)

OpenStudy (anonymous):

No problem!

OpenStudy (anonymous):

theres another one can u help me?

OpenStudy (anonymous):

Peg works 2 1/2 hours each day on 4 days a week sewing covers. At $11.25 per hour, how much does she earn each week?

OpenStudy (anonymous):

hours x wage ($ per hour) = earning. so plug the values given in the question into the equation

OpenStudy (anonymous):

how do u do the 2 1/2 thoee?

OpenStudy (anonymous):

\[2\frac{ 1 }{ 2 }\]= 2.5

OpenStudy (anonymous):

but it doesnt give me an answer thoe?

OpenStudy (anonymous):

to find how many hours Peg worked that week, take the 2.5 and multiply it by the 4 days she worked

OpenStudy (anonymous):

thats 10 , riqht?

OpenStudy (anonymous):

correct now, take the number of hours she worked (10) and multiply it by the amount of money she makes an hour.

OpenStudy (anonymous):

so it will be $112.50

OpenStudy (anonymous):

Correct!

OpenStudy (anonymous):

oo, :)

OpenStudy (anonymous):

is this question the same as u just helped me w ? Kyle works in the stock department of a discount toy store for 8 hours each on Saturdays and Sundays. One weekend he earned $76. How much did he earn per hour?

OpenStudy (anonymous):

Almost, you're just working backwards, but it's the same concept.

OpenStudy (anonymous):

what u mean workinq backwards?

OpenStudy (anonymous):

division

OpenStudy (anonymous):

earning / hours = wage

OpenStudy (anonymous):

i got 9.50?

OpenStudy (anonymous):

He works 8 hours 2 different days, so you would divide 76 by 16, not 8.

OpenStudy (anonymous):

oo iic :)

OpenStudy (anonymous):

good!

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

what does units have to do w money? this is my question.! 102 units at $8.72 each

OpenStudy (anonymous):

units is referring to anything that you can buy. it's a variable, like x or y. it's could be a pack of socks, a mcdonalds meal, ect in real life.

OpenStudy (anonymous):

so i divide those 2 numbers?

OpenStudy (anonymous):

1 unit costs $8.72, two units cost $17.44, three units cost $26.16 and so on. division causes numbers to get smaller, and these numbers aren't getting smaller.

OpenStudy (anonymous):

riqht so do ii keep on goinq till i get to 102??

OpenStudy (anonymous):

multiplication would be much faster lol. 8.72 x 102

OpenStudy (anonymous):

would ii do the same w tips? $2.95 per hour; 30 hours; $295.25 tips

OpenStudy (anonymous):

remember our formula: wage x hours = earnings. because a waitress would not receive $295.25 in tips every hour, you have to add tips in separately. with tips, our formula would be (wage x hours) + tips = earnings

OpenStudy (anonymous):

so the answer would be 383.75?

OpenStudy (anonymous):

correct!

OpenStudy (anonymous):

oo okaii thanks :)

OpenStudy (anonymous):

is this the same? Hours: 20; Hourly rate: $3.25; Weekly tips: $115.75, $94.80, $73.15, $62.50

OpenStudy (anonymous):

yep! add all of the tips together and then plug them into the formula

OpenStudy (anonymous):

so ii multiply then add the tips riqht?

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

so ii did everythinq a ii got a huqee number ha

OpenStudy (anonymous):

what did you put into your calculator?

OpenStudy (anonymous):

i put 20 x 3.25=65 65+115.75=180.75+94.80=275.55+73.15+73290.55+62.50=73353.05

OpenStudy (anonymous):

then i did it aqain and ii got 411.20

OpenStudy (anonymous):

411.20 is correct. "73290.55" is a typo that threw your number waaay off haha

OpenStudy (anonymous):

haha yeaa it did , but the problem is that its not on my answers i onlly have 138.15, 151.55, 148.00

OpenStudy (anonymous):

ups forgot to tell u , its lookinq for the average weekly earninqs

OpenStudy (anonymous):

ah! okay, do you know how to find the average of a set of numbers?

OpenStudy (anonymous):

um i think, dont u add all them then divide by how many sets der is?

OpenStudy (anonymous):

yup! so find the average of the tips and then add it to the $65 she makes per week.

OpenStudy (anonymous):

how u get 65?

OpenStudy (anonymous):

nvm i got it

OpenStudy (anonymous):

look at this?

OpenStudy (anonymous):

Unit rate: $0.76; Units per week: 502, 398, 433, 480

OpenStudy (anonymous):

are you finding the average again?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay, average up the units/week and then multiply that number by the unit rate

OpenStudy (anonymous):

so its 344.47

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

well thanks so much for all ur help, ii have to head to ma bed cuz im really tierd , but thanks once aqain for evertyhinqq :), nitee, & sweet dreams :) hope to talk to u tomorrow if ii need help :D

OpenStudy (anonymous):

haha anytime! sleep well! c:

OpenStudy (anonymous):

Youu Too, THanks :)